OpenStudy (anonymous):

the perimeter of a retangular sign is not to exceed 50ft. the length is to be twice the width. what widths will meet these conditions?

OpenStudy (anonymous):

do you know the formula for the perimeter?

OpenStudy (anonymous):

....?

OpenStudy (anonymous):

doesnt sy that is all the question that it gives

OpenStudy (anonymous):

but yoy should know that P=2l+2w

OpenStudy (anonymous):

i havent been in class long and dont know anything about algebra lol

OpenStudy (anonymous):

im 35 trying to go back to school

OpenStudy (anonymous):

ok:) if l is the length of the rectangle w is its width then P=2l+2w the length is twice the width=>l=2w then the perimeter becomes:P=2(2w)+2w=6w i hope you're following so far?

OpenStudy (anonymous):

yes i am thanks

OpenStudy (anonymous):

ok,it says the perimeter is not to exceed 50 which means P<=50 ( less than or equal to) =>6w<=50 =>w<=50/6 w<=25/3 this means all widths which are less than or equal to 25/3 ft are accepted

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

so thats my answer?

OpenStudy (anonymous):

yes its not only one answer cause its an inequality:)

OpenStudy (anonymous):

oh ok i understand now .. just wish i could understand how you did all that lol im going to write it down andd see if i can do it lol...thank you

OpenStudy (anonymous):

ok,if you have any questions,just ask:)

OpenStudy (anonymous):

ok thanks i will

OpenStudy (anonymous):

convert to a percent 0.00763 isnt that move dec 2 times to the right?

OpenStudy (anonymous):

or the left?

OpenStudy (anonymous):

to the right,or multiply by hundred

OpenStudy (anonymous):

763%

OpenStudy (anonymous):

0.763%

OpenStudy (anonymous):

one by one...% to decimal 2 places to the left

OpenStudy (anonymous):

k thanks